Compare the energy (by ratio!) of one photon of visible light to one photon of an X ray. Select a particular (nominal) wavelength for each.
Answer #1

E=h

where h is plancks constant and is frequency of photon

for visible light range is 405*1012 to 790*1012 Hz

for visible light let us take a nominal frequency of 600*1012 Hz

For X-ray range is 3×1016 Hz to 3×1019 Hz

for x-ray let us take a nominal frequency of 3*1018Hz

Energy ratio=frequency ratio=(600*1012)/(3*1019)=2×10−5